(D^2-2D+3)y=X^2-1


1
Expert's answer
2021-05-24T15:20:46-0400

Let us solve the differential equation "(D^2-2D+3)y=x^2-1". For this let us solve the characteristic equation of the homogeneous equation "(D^2-2D+3)y=0":


"k^2-2k+3=0"


"(k-1)^2+2=0"


"(k-1)^2=-2"


"k_1=1+i\sqrt{2}" and "k_2=1-i\sqrt{2}."


It follows that the general solution of the equation "(D^2-2D+3)y=x^2-1" is of the form "y=e^x(C_1\cos(\sqrt{2}x)+ C_2\sin(\sqrt{2}x))+y_p," where "y_p=Ax^2+Bx+C." It follows that "y_p'=2Ax+B" and "y_p''=2A." Therefore, "2A-2(2Ax+B)+3(Ax^2+Bx+C)=x^2-1." We conclude that "3Ax^2+(3B-4A)x+(2A-2B+3C)=x^2-1." Consequently, "3A=1,\ 3B-4A=0,\ 2A-2B+3C=-1." It follows that "A=\frac{1}{3},\ B=\frac{4}{9},\ C=\frac{1}{3}(-1-\frac{2}{3}+\frac{8}{9})=-\frac{7}{27}."


We conclude that the general solution of the differential equation "(D^2-2D+3)y=x^2-1" is


"y=e^x(C_1\cos(\sqrt{2}x)+ C_2\sin(\sqrt{2}x))+\frac{1}{3}x^2+\frac{4}{9}x-\frac{7}{27}."